OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2). Answer y = 5/2 x - 2 y = -5/2 x + 8 Y = 2/5 x - 6 y = -2/5 x + 4

OpenStudy (anonymous):

slope = 8 +2 /4 = 10/4 = 5/2 y = mx + c -2 = 0 + c c = -2 equation = y = 5/2x - 2

OpenStudy (anonymous):

slope formula : slope(m) = (y2 - y1) / (x2 - x1) (4,8) -- x1 = 4 and y1 = 8 (0,-2) -- x2 = 0 and y 2 = -2 now les sub in our stuff and find the slope... m = (-2 - 8) / (0 - 4) m = -10/-4 m = 5/2 now we will use point slope formula : y - y1 = m(x - x1) (4,8) -- x1 = 4 and y1 = 8 slope(m) = 5/2 y - 8 = 5/2(x - 4) y - 8 = 5/2x - 10 y = 5/2x - 10 + 8 y = 5/2x - 2 I agree with joel :)

OpenStudy (anonymous):

WOAH , WOAH @kelliegirl33

OpenStudy (anonymous):

:)